OpenStudy (anonymous):

Algebra II What is the simplified form of the quantity 1 over x minus 1 over y all over 1 over x plus 1 over y?

OpenStudy (faiqraees):

\[\large\rm \frac{\frac{1}{x}-\frac{1}{y} }{\frac{1}{x}+\frac{1}{y} } \]\[\large\rm \frac{\frac{y-x}{xy} }{\frac{y+x}{xy}} \]\[\large\rm \frac{y-x}{y+x} \]